OpenStudy (owlet):

Find the absolute ,maximum and minimum of \(\sf f(x,y)= 4 x y^2-x^2y^2-xy^3\) on the close triangular region in the xy-plane with vertices (0,0), (0,6), (6,0).

OpenStudy (owlet):



OpenStudy (owlet):

\(\sf fx= 4y^2-2xy^2-y^3=0\\y^2(4-2x-y)=0\\ .\\fy=8yx-2x^2y-3y^2x=0\\xy(8-2x-3y)=0\) From that, the points I found are: (0,0,0) and (1,2,4)

OpenStudy (owlet):

Now from the boundaries L1: x=0, 0<= y<=6, I'll get (0,0,0) L2: y=0, 0<=x<= 6, I'll get (0,0,0) L3: y=-x+6 I sub it in the original function and then I evaluated its derivative and found x=-6 and x=6 as roots, since x=-6, is not in the boundary, I removed it and used x=6 instead. When x=6, y=0...so I will have point (6,0,0) So my absolute max will be (1,2,4) and my min are (0,0,0) and (6,0,0)...

OpenStudy (owlet):

However the right answer for the minimum is (2,4,-64)

ganeshie8 (ganeshie8):

The boundary of the region are not points. They are the 3 sides of the triangle : x+y = 6 x = 0 y = 0